OpenStudy (anonymous):

If 2az-5z=2 then z=?

Directrix (directrix):

2az-5z=2 z( 2a - 5) = 2 z = 2/ (2a -5)

OpenStudy (anonymous):

can u explain what you did please

Directrix (directrix):

1. 2az-5z=2 ---> the problem 2. To solve for z, then the z terms have to be collected on one side of the equation. They are already collected on one side. 3. Factor out the z term. z ( 2a - 5) = 2 4. Think of this as a coefficient times a variable. as in (2a - 5) * z = 2 [Note: If we had 7Z = 2, the answer would be z = 2/7.] We will do the same for (2a - 5) * z = 2. 5. z = 2 / (2a - 5) Note that solving for z means there are no z terms in the solution for z.